The measure of one angle of a triangle is five times that of a second angle, and the measure of an exterior angle at the third vertex is 120 degrees. What is the mesure of each angle of the triangle?
Answer by checkley71(8403)

first angle =x, the second is 5x & the third angle is 180-120=60
so x+5x+60=180
6x=180-60
6x=120
x=120/6
x=20 the 3 angles are 20, 100 & 60.
proof
20+5*20+60=180
20+100+60=180
180=180